OpenStudy (anonymous):

Help please! sin(90-theta)csc theta / cot theta =1

OpenStudy (anonymous):

Use trig identities to transform the left side to the right. Here's the equation in better form...\[\frac{ \sin(90-\theta)\csc \theta }{ \cot \theta }=1 \]

OpenStudy (anonymous):

sin(90 - theta) = sin(90)cos(90) + cos(theta)sin(theta) = cos(theta)sin(theta)

OpenStudy (anonymous):

I meant a minus sign not plus. -cos(theta)sin(theta)

OpenStudy (anonymous):

\[\frac{ (\cos(\Theta)\sin(90)(\sin(\Theta) }{ \cos(\Theta)\sin(\theta) }\]

OpenStudy (anonymous):

simplify.

OpenStudy (anonymous):

So... (cos(theta)(sin(theta)/(cos(theta)(sin(theta) and then it all equals 1?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks! :D